SOAR Telescope resumes science operations

Submitted by jelias on Sun, 2020-09-27 22:57

The SOAR Telescope was closed due to the coronavirus pandemic on March 16. (See previous article as well as NOIRLab announcements and updates here.)

Following a successful review of restart plans for AURA's Chilean operations, held on September 11, limited resumption of science operations was authorized to start on September 28.

The first steps in this process were careful inspections and critical maintenance of the telescope, dome and auxiliary facilities, with paralle work pumping and cooling the science instruments. This work is being done with appropriate safety precautions in place.

Observers should understand that operations will be occurring with significant requirements on social distancing and limitations on group activities. As a result, we may not be able to support complex set-ups, and responses to failures may be slower than usual. In particular, there may be instances where telescope time is lost because of delayed repairs. We appreciate observers' understanding in this regard.

Questions may be directed to the undersigned, your staff contact, or your institutional telescope time coordinator.

As stated above, SOAR restart operations, as well as parallel activities at the other AURA facilities, are following a detailed plan that was fully reviewed and approved in mid-September. Staff have been briefed on the plan and trained on associated safety procedures and precautions. We believe that adherence to the plan will allow us to re-open and continue to operate while protecting our staff.

We have been contacting scheduled observers, based on an amended schedule.

November 14 status update (final)

The telescope and Goodman spectrograph were ready for the first night of engineering October 5, however poor weather intervened and we were not able to open until the following night. This was sufficient to check out the telescope and instrument (some further checks were performed the following night); science operations resumed October 7 and are continuing.

Subsequently, all of our instruments have been prepared for use, and all instruments scheduled for the remainder of the semester have been checked on sky, now including science observations. Aside from Goodman, which has seen regular use since October 7, observations have been performed with SIFS, HRCam, SAM, Spartan and TripleSpec, plus a complete day-time check-out with SOI. Auxiliary capabilities (guiders, ADC, calibration WF sensor) have all been brought into operation as well.

The NOIRLab main web page has a press release on lab-wide operations restart, including some SOAR pictures. Our successful restart is the result of hard work, performed safely, by the SOAR staff and supporting members of CTIO Engineering and Technical Services, as well as NOIRLab Central Operations Services staff.

This should be considered the final update to this web page. We will likely post a new item describing plans for the upcoming semester (2021A) in the next few weeks.

- Jay Elias, September 27, October 2, October 12, October 25, November 14

Updated on June 2, 2022, 6:58 am